Doon Iowa Derailment

Doon, IA - Region VII

 

Site Contact:

Melinda Luetke, Heath Smith
OSC

Luetke.Melinda@epa.gov, Smith.Heath@epa.gov

Doon, IA
Doon, IA 51235
response.epa.gov/DoonIADerailment



NRC#: 1216081

At 0624 HRS (CST) June 22, 2018 the EPA Region 7 Spill Line was contacted by the NRC to report a BNSF train derailment in Doon, IA. The derailment is located in flood impacted waters of the Rock River. Derailed cars contain heavy crude/diluent mix, material has been released and is visible in flood waters. IDNR is providing updates to status and has dispatched two responders to the incident. Two EPA OSCs have been mobilized to the site along with START contractors.

Waters of the Rock River flow to the Big Sioux River in South Dakota and continue to the Missouri River north of Sioux City, IA. EPA is notifying downstream stakeholders including Sioux City Emergency Management, Sioux City Drinking Water, Nebraska Department Environmental Quality, and the South Dakota Department of Environment and Natural Resources. Tribes of Omaha and Winnebago have been notified of release. Local responders have reported that well fields for Rock Valley, Iowa, have been shut down. EPA Drinking water program coordinated to provide notifications to IDNR water supply staff and the City of Rock Valley Rural water district. Osceola County and the city of Sheldon were also under advisory.

The Regional Environmental Protection Specialist at DOI's Office of Environmental Policy and Compliance provided information concerning USFWS resources in the area, they specified that the endangered Topeka Shiner is present in the area. USFWS South Dakota field office & their spill coordinator in IA continue to coordinate for the spill. EPA Region 8 has contacted the R7 Spill line to coordinate and assist as needed. The Big Sioux River is also at flood stage and is located in South Dakota part of EPA Region 8 jurisdiction.

EPA OSCs continue updates to R7 REOC as personnel transit to site. At 1215 hours on 6/22/2018, BNSF reported 31 cars are derailed and are rolled over, 2 cars have derailed and are upright. BNSF states that petroleum materials released are heavy crude/diluent mix, SDS is being forwarded to EPA REOC. Initial contract personnel are arriving onsite, air monitoring contractors CTEH are expected at 1330 hours on 6/22/2018 and additional response contract personnel are expected later on the same day.

BNSF provided a general strategy for response which includes attempts to contain source at area of derailment and deploying downstream personnel to determine additional points for containment. Flood waters will complicate response activities and will continue to be significant while rain is forecasted through early next week. Further updates will continue as response develops.




Resources

Notices

  • None for this site

Images

  • None for this site

Documents

  • None for this site

Contacts

  • None for this site

Links

  • None for this site